Please help me !!!
Use an equiprobable sample space to commute the probability of getting exactly 3 heads in 6 tosses of a fair coin.

Answer by richard1234(7193)

P(3/6 heads) = (6C3)/2^6 = 20/64 = 5/16

Here, nCr denotes the binomial coefficient or "n choose r."